Krieger 205 Seating Layout File kr205, Revision D, jmb 3/99 Johns Hopkins University

There are 165 seats, including 9 left-handed seats. With alternate seating, there is space for 86 people, provided at least 5 of them are left-handed. Every third seat is labeled. There is no row I. The seats are not padded.

⋆ Left-handed seats are available at the left end of most rows. † Seats A1 and B1 are designated handicapped-accessible.

L 15 14 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ L

K 18 17 16 15 14 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ K

J 18 17 16 15 14 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ J

H 18 17 16 15 14 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ H

G 18 17 16 15 14 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ G

F 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ F

E 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ E

D 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ D

C 13 12 11 10 9 8 7 6 5 4 3 2 1⋆ C

B 13 12 11 10 9 8 7 6 5 4 3 2 1† B

A 13 12 11 10 9 8 7 6 5 4 3 2 1† A

BENCH DOOR ↓